OpenStudy (anonymous):

If AX:XB as AB:AC, find the missing length. AX = 5, XB = 4, AC = 6, find AB.

OpenStudy (mathmagician):

5/4=AB/6 AB=(6*5)/4=7.5

OpenStudy (mertsj):

\[\frac{5}{4}=\frac{x}{6}\]

OpenStudy (anonymous):

15/2-7.5